OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations? x2 + y2 = 36 y = x - 6 A. 2 and 0 B. 6 and 2 C. 6 and -2 D. 6 and 0

OpenStudy (anonymous):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Is this what the eqautions are? \[x^2+y^2 = 36\] and \[y=x-6\]

OpenStudy (anonymous):

yes

OpenStudy (yanasidlinskiy):

the second equation, pick a number for x and the results will be for why, and from their. Keep doing that until u have like 3 ordered pairs and draw a line for it . y = x - 6 and then for the second one, all you have to do is, change the equation and solve for y. x2 + y2 = 36 change it to y^2 =36 + x^2

OpenStudy (yanasidlinskiy):

Go it? I have to go now.

OpenStudy (anonymous):

substitute the second equation into the first and solve for x. There should be 2 values of x. Then substitute the 2 values of x in turn into the second equation to get the 2 corresponding values of y.

OpenStudy (anonymous):

You ok with it?

OpenStudy (anonymous):

how would i substitute the second equation into the first

OpenStudy (yanasidlinskiy):

x^2 + y^2 = 36 square everything x + y = 6 x + y = 6 y = x - 6 x+ (x-6) = 6 2x - 6 = 6 2x = 12 x = 6 substitute that into y = x- 6 to get the value for y :)

OpenStudy (anonymous):

sorry only asking for the x co-ordinate so no need to substitute back into the second

OpenStudy (anonymous):

may I continue @bribri50?

OpenStudy (anonymous):

you may

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so the answer is either b, c, or d

OpenStudy (anonymous):

\[x^2+y^2=36\]\[y=x-6\]\[x^2+(x-6)^2=36\]

OpenStudy (anonymous):

\[x^2+x^2-12x+36=36\]\[2x^2-12x=0\]

OpenStudy (yanasidlinskiy):

Oh, shhoot. Sorry for the incorrections. I have an explanation.. First note that we take y = x - 6 and substiutute that into y in the first equation, So the FIRST equation becomes x^2 + (x-6)^2 = 36 Simplify the above, you get x^2 + x^2 - 12x + 36 = 36 2x^2 - 12 x = 0 2x(x - 6) = 0 x = 0 or x = 6 If x = 0, y = -6; if x = 6, y = 0 Final solutions: x = 0 and y = -6 AND x = 6 and y= 0

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

\[x^2-6x=0\]\[x(x-6)=0\] \[x=0 or x=6\]

OpenStudy (anonymous):

@YanaSidlinskiy got it before me.

OpenStudy (yanasidlinskiy):

Thanx for noticing my mistake. Hahahaha:D I gave you a medal though:D